OpenStudy (anonymous):

the area of a triangle is 6.75 m squared. if the base of the triangle is 3 m, what is the height of the triangle?

OpenStudy (debbieg):

Work it out backwards from the formula for area of a triangle: area = (0.5)bh

OpenStudy (debbieg):

Solve for h, plug in your known area and base, and you'll get height.

OpenStudy (anonymous):

So, would my equation be 3*h=6.75?

OpenStudy (debbieg):

\[\large A=\frac{ 1 }{ 2 }bh\] So plugging in your known values, you get: \[\large 6.75=\frac{ 1 }{ 2 }3h\] so: \[\large 6.75=\frac{ 3 }{ 2 }h\] So to solve for h, what do you need to do?

OpenStudy (anonymous):

would I need to divide?

OpenStudy (debbieg):

Well, the easiest way would be to multiply both sides by the reciprocal of the coefficient (3/2), right? That would leave just h on the RHS and the solution for h on the LHS

OpenStudy (anonymous):

OH. okay, so I would multiply the reciprocal of 3/2, which would be 2/3 correct?